### v2

# Noncritical Conformal Gravity and Four-Dimensional Liouville Theory

Abstract
We study the quantum aspects of the conformal gravity in four dimensions, specifically addressing a known discrepancy in beta functions between general quadratic curvature theories and conformal gravity, which corresponds to two scalar degrees of freedom. We demonstrate that this mismatch is resolved by carefully introducing gauge-fixing and ghost terms via the BRST symmetry, which effectively adds the two scalar modes. Drawing lessons from two-dimensional quantum gravity and Liouville theory, we proceed to integrate the four-dimensional trace anomaly to derive a consistent Liouville action, which is given by a free-field action for the conformal mode with a consistent conformal anomaly. Finally we give the condition that the BRST transformation is anomaly free.
1 Introduction
In this paper we study the quantization of conformal gravity in four dimensions. The conformal sector is described by the conformal factor $e^{\phi(x)}$ in spacetime metric of the form
$g_{\mu\nu}(x)=e^{\phi(x)}{\hat{g}}_{\mu\nu}(x),$ (1.1)
where ${\hat{g}}_{\mu\nu}$ is a family of reference metric representing the degrees of freedom apart from the conformal mode. Before going into this subject, let us first briefly summarize a puzzle in the beta functions for quadratic curvature theory, which helps us to better understand the theory.
It has been known for some time that there is a subtlety in the beta functions of the general quadratic curvature theory
$S=\int d^{4}x\sqrt{g}\,\left(\frac{1}{2\lambda}C_{\mu\nu\alpha\beta}^{2}+\frac{1}{\xi}R^{2}-\frac{1}{\rho}R_{\rm GB}^{2}\right),$ (1.2)
and the conformal gravity
$S=\int d^{4}x\sqrt{g}\,\left(\frac{1}{2\lambda}C_{\mu\nu\alpha\beta}^{2}-\frac{1}{\rho}R_{\rm GB}^{2}\right),$ (1.3)
where $\lambda,\xi$ and $\rho$ are dimensionless couplings, and $C_{\mu\nu\alpha\beta}$ is the Weyl tensor which is defined such that its trace is zero. Its square is given as
$C_{\mu\nu\lambda\sigma}^{2}=R_{\mu\nu\lambda\sigma}^{2}-2R_{\mu\nu}^{2}+\frac{1}{3}R^{2},$ (1.4)
and $R_{\rm GB}^{2}$ is the Gauss-Bonnet (GB) term
$R_{\rm GB}^{2}=R_{\mu\nu\alpha\beta}^{2}-4R_{\mu\nu}^{2}+R^{2},$ (1.5)
Tbe beta functions in this system may be calculated by the functional renormalization group equation
$k\frac{d}{dk}\Gamma_{k}=\frac{1}{2}{\rm Tr}\left(\Gamma_{k}^{(2)}+R_{k}\right)^{-1}k\frac{d}{dk}R_{k},$ (1.6)
where $\Gamma_{k}$ is the effective average action, $k$ is the cutoff, $R_{k}$ is a cutoff function suppressing the contribution of the modes below the momentum scale $k$ to $\Gamma_{k}$, and $\Gamma_{k}^{(2)}$ is the second variation of the effective average action. The right-hand side of (1.6) gives the anomaly coeffectients $a,b$, anc $c$:
$\int d^{4}x\sqrt{g}\frac{1}{(4\pi)^{2}}\left(c\,C_{\mu\nu\lambda\sigma}^{2}+bR^{2}-a\,R_{\rm GB}^{2}\right).$ (1.7)
in terms of which the beta functions are given as
$\beta_{\lambda}$ $=$ $-\frac{2c}{(4\pi)^{2}}\lambda^{2},$
$\beta_{\xi}$ $=$ $-\frac{b}{(4\pi)^{2}}\xi^{2},$
$\beta_{\rho}$ $=$ $-\frac{a}{(4\pi)^{2}}\rho^{2}.$ (1.8)
The beta functions for the general quadratic curvature theory have been calculated in [1, 2, 3, 4], while those for the conformal gravity have been calculated in [1, 5, 6]. Naively it is expected that the beta functions for the conformal gravity are reproduced if we take the limit $\xi\to\infty$ [8, 7]. Surprizingly enough, it is found that they do not agree. Rather there are discrepancy corresponding to precisely two scalar degrees of freedom. This may arise because the conformal gravity, to which the general quadratic theory appears to reduce in the $\xi\to\infty$ limit, has additional gauge symmetry under the conformal transformation; in this limit, the conformal symmetry is recovered, so without fixing it, the theories may not coincide. The problem is recently understood by making partial gauge fixing in the conformal gravity [9]. We then discuss the resolution, and find the formulation can be used to derive the consistent quantum conformal gravity with four-dimensional Liouville theory.
This paper is organized as follows. In Sect. 2, based on the BRS transformation for the conformal invariance, we fix the degrees of freedom of the conformal modes and evaluate the beta functions. Indeed, we can confirm that the discrepancy mentioned above can be explained as the contribution of the ghost and anti-ghost fields for this gauge fixing. Section 3 reviews the derivation of the Liouville action in two dimensions and the conditions for the correct quantization of the conformal modes. Using this as a reference, Sect. 4 derives the Liouville action for the four-dimensional case and provides consistent quantization conditions for the conformal modes. In the two-dimensional case, the quantization condition was that the total central charge involving the Liouville field vanishes. Similarly, for the four-dimensional case, we can confirm that this condition is replaced by the requirement that the total beta function involving the Liouville field and the ghosts of the conformal transformation vanishes. We show that this uniquely leads to a free field action for the conformal mode. We also contrast this approach with the widely-used Riegert proposal [10] for the four-dimensional Liouville action, providing a more robust, consistent formulation derived through the quantization procedure. We summarize our results in Sect. 5. Some necessary formulae and discussions are relegated to the appendices. In Appendix A, we display the transformation property of the curvaturs under the conformal transformation. In Appendix B, we discuss the transformation properties of scalar field under the conformal transformation. Finally in Appendix C, we summarize the consistency condition for the trace anomaly.
2 Conformal anomaly from the conformal gravity
Let us consider the action for the conformal gravity (1.3). Since this has the conformal symmetry under the Weyl transformation $g_{\mu\nu}\to e^{\sigma}g_{\mu\nu}$, to quantize the system, we have to gauge fix it. To identify the gauge fixing and the corresponding Faddeev-Popov ghost terms, we use the BRST symmetry. The infinitesimal conformal or more precisely Weyl transformation is given as
$\delta g_{\mu\nu}=\epsilon g_{\mu\nu},$ (2.1)
where $\epsilon$ is an infinitesimal transformation parameter. This is promoted to the BRST transformation
$\delta_{B}g_{\mu\nu}=\delta\lambda\,cg_{\mu\nu},$ (2.2)
where $\delta\lambda$ is an anticommuting parameter and $c$ is a ghost field for the conformal transformation. [^1] This ghost $c$ and the following antighost $b$ should not be confused with the $a,b,c$ coefficients in Eq. (1.7). The BRST transformation of the ghost is determined by the requirement that the BRST transformation (2.2) be nilpotent. This gives
$\delta_{B}c=0.$ (2.3)
Finally the BRST transformations of the antighost $b$ and the associated auxiliary field $B$ would be naively defined as
$\delta_{B}b\stackrel{{\scriptstyle?}}{{=}}\delta\lambda B,\qquad\delta_{B}B\stackrel{{\scriptstyle?}}{{=}}0.$ (2.4)
However we expect that the ghosts $(b,c)$ have the conformal weight (1,0)(-1,0). The transformation (2.4) does not take into account the conformal weight of the antighost $b$. It turns out [9] that we can incorporate the conformal weight by the following modifications
$\delta_{B}b=\delta\lambda(-cb+B),\qquad\delta_{B}B=-\delta\lambda cB,$ (2.5)
which are still nilpotent. Under this transformation, we find from the result in Appendix A that the determinant of the metric and the scalar curvature transform as
$\delta_{B}\sqrt{g}=2\delta\lambda c\sqrt{g},\qquad\delta_{B}R=-\delta\lambda(cR+3\square c).$ (2.6)
We introduce $R$ as a partial gauge fixing function of the conformal symmetry:
$f=R.$ (2.7)
Conformal invariant quantities can be computed integrating over an arbitrary smearing of the gauge condition around $f=0$. The gauge fixing term and the Faddeev-Popov ghost term may then be written as [11, 12]
${\cal L}_{\rm GF+FP}\hskip-5.69054pt$ $=$ $\hskip-5.69054pt\delta_{B}\left[\sqrt{g}b\left(R-\frac{\alpha}{2}B\right)\right]/\delta\lambda$ (2.8)
$=$ $\hskip-5.69054pt\sqrt{g}\left[2cb\left(R-\frac{\alpha}{2}B\right)+(-cb+B)\left(R-\frac{\alpha}{2}B\right)+b(cR+3\square c)-\frac{\alpha}{2}bcB\right],\qquad$
where $\alpha$ is a gauge parameter. This gives the action
$S_{\rm GF+FP}=\int d^{4}x\sqrt{g}\left[-\frac{\alpha}{2}\left(B-\frac{1}{\alpha}R\right)^{2}+\frac{1}{2\alpha}R^{2}+3b\square c\right],$ (2.9)
where anticommuting property of the ghosts is used so that bc=cbbc=-cb. [^2] If we use the transformation (2.4), the same procedure gives $S_{\rm GF+FP}=\int d^{4}x\sqrt{g}\left[-\frac{\alpha}{2}\left(B+cb-\frac{1}{\alpha}R\right)^{2}+\frac{1}{2\alpha}R^{2}+3b\square c\right].$ So practically both transformations give the same result. The original conformal transformation is broken by the gauge fixing term, leaving only the global BRST invariance. Thus we may adopt transformations for the antighost and auxiliary fields different from those suggested by the conformal transformation as long as we consider the proper transformation for the other physical fields. It is interesting that there is no $Rbc$ term in the action (2.9) even though $b$ and $c$ are (fermionic) scalar fields, in contrast to usual bosonic scalar fields (see Appendix B).
[…]
After the gauge fixing of the reparametrization, by use of the heat kernel expansion, the rhs of this equation was calculated as [1, 2, 3, 4]
$\int d^{4}x\sqrt{g}\frac{1}{(4\pi)^{2}}\left[\frac{133}{20}C_{\mu\nu\lambda\sigma}^{2}+\frac{5(72\lambda^{2}-36\lambda\xi+\xi^{2})}{36\,\xi^{2}}R^{2}-\frac{196}{45}R_{\rm GB}^{2}\right],$ (2.10)
where only the $b_{4}$ coefficients are kept. The beta functions for the quadratic terms are then
$\beta_{\lambda}^{\,\rm g}$ $=$ $-\frac{1}{(4\pi)^{2}}\frac{133}{10}\lambda^{2},$
$\beta_{\xi}^{\,\rm g}$ $=$ $-\frac{1}{(4\pi)^{2}}\frac{5}{36}(72\lambda^{2}-36\lambda\xi+\xi^{2}),$ (2.11)
$\beta_{\rho}^{\,\rm g}$ $=$ $-\frac{1}{(4\pi)^{2}}\frac{196}{45}\rho^{2}.$
If we consider the limit of $\xi\to\infty$ in Eq. (2.10), we expect that this would give the result for the conformal gravity without $R^{2}$ term, and the beta function for $\xi$ is absent. However, the beta functions for the conformal gravity was calculated in several works [1, 5, 6], and the result is given as
$\beta_{\lambda}^{\,\rm conf}$ $=$ $-\frac{1}{(4\pi)^{2}}\frac{199}{15}\lambda^{2},$
$\beta_{\rho}^{\rm conf}$ $=$ $-\frac{1}{(4\pi)^{2}}\frac{87}{20}\rho^{2},$ (2.12)
which are different from the limit of $\xi\to\infty$ in (2.11). This had been a mystery for some time.
The above result is then written in terms of $a,b,c$ coefficients in (1.7) as
$c^{\,\rm g}=\frac{133}{20},\quad$ $a^{\,\rm g}=\frac{196}{45},$ (2.13)
$c^{\,\rm conf}=\frac{199}{30},\quad$ $a^{\,\rm conf}=\frac{87}{20},$ (2.14)
Their difference is
$c^{\,\rm g}-c^{\,\rm conf}=\frac{1}{60},$
$a^{\,\rm g}-a^{\,\rm conf}=\frac{1}{180}.$ (2.15)
On the other hand, it is known that the contribution of the matter is
$c^{\rm m}=\frac{1}{120}(N_{S}+6N_{F}+12N_{V}),$
$a^{\rm m}=\frac{1}{360}(N_{S}+11N_{F}+62N_{V}),$ (2.16)
where $N_{S},N_{F}$ and $N_{V}$ are the number of scalar, fermion and vector fields, respectively. It is then clear that there is mismatch in the coefficients (2.15) precisely by the contribution of two scalar degrees of freedom [1, 8, 9, 7]. Interesting enough, the above formulation with gauge fixing function (2.7) beautifully explains the mismatch. We have seen that the gauge fixing procedure introduces precisely two scalar (fermionic) modes $b$ and $c$. This is the reason why there is an apparent mismatch of two scalar degrees of freedom. The gauge fixing and ghost terms (2.9) actually decouple from the physical space. This is consistent with the fact that the beta function for $\lambda$ is independent of $\xi$.
However, the gauge-fixed theory constructed as above cannot necessarily be regarded as a conformally invariant gravitational theory. This is because the conformal invariance is broken unless the total beta functions are zero (see (2.12)):
$\frac{199}{30}+c^{\rm m}=0,$
$\frac{87}{20}+a^{\rm m}=0.$ (2.17)
Generically, except in special cases satisfying the above conditions, even if a system including gravity is classically conformally invariant, quantization breaks the symmetry due to the anomaly. In this case, the conformal modes do not appear in the classical action but do appear in the effective action. Therefore, the path integral for the conformal modes requires careful consideration. This is a four-dimensional analogue of noncritical strings in two dimensions and may be called noncritical conformal gravity.
In two dimensions, this problem has been resolved by Liouville theory. To study this in four dimensions, let us start by reviewing the two-dimensional situation.
3 Lessons from two-dimensional quantum gravity for conformal gravity – Integrating the trace anomaly
The action of two-dimensional gravity is
$S_{2}=\int d^{2}\xi\sqrt{g}\left(R+\mu\right),$ (3.1)
where we have set the coefficient of the Einstein term to be 1 and $\mu$ is the cosmological constant. We use the Eucleadean metric. The Einstein term in two dimensions gives just topological number
$\int d^{2}\xi\sqrt{g}R=4\pi\chi=\mbox{const.}$ (3.2)
The energy-momentum tensor is given by the partition function $Z[g]$ as
$\langle T_{\mu\nu}\rangle$ $=$ $-\frac{2}{\sqrt{g}}\frac{\delta\log Z[g]}{\delta g^{\mu\nu}}.$ (3.3)
This implies that
$\delta\log Z[g]=-\int d^{2}\xi\frac{1}{2}\sqrt{g}\,\delta g^{\mu\nu}T_{\mu\nu}.$ (3.4)
When the conformal matter is coupled, there appears trace anomaly. In such a case, under the scale transformation $\delta g^{\mu\nu}=-g^{\mu\nu}\delta\phi$, we have the relation
$\delta(\log Z[g]):=\log(Z[ge^{\delta\phi}])-\log(Z[g])=\int d^{2}\xi\,\frac{1}{2}\,\sqrt{g}\,\delta\phi\,T^{\mu}_{\mu}$ (3.5)
The trace anomaly is given by
$\langle T^{\mu}_{\mu}\rangle=\frac{c}{24\pi}R,$ (3.6)
where $c$ is called the central charge.
Write the partition function as
$Z=e^{-F},$ (3.7)
and we have
$\delta F=F[ge^{\delta\phi}]-F[g]=-\frac{c}{48\pi}\int d^{2}\xi\sqrt{g}R\delta\phi.$ (3.8)
Now we write $\delta\phi=\phi dt$, and Eq. (3.8) is
$dF=-\frac{c}{48\pi}\int d^{2}\xi\sqrt{g(t)}R(t)\phi dt,$ (3.9)
where $R(t)$ is the curvature for $g(t)=ge^{t\sigma}$. Using the formula (A.2) and integrating (3.9) from $t=0$ to $t=1$, we obtain
$F[ge^{\phi}]-F[g]$ $=-\int_{0}^{1}\frac{c}{48\pi}\int d^{2}\xi\sqrt{g}(R-t\square\phi)\phi dt$
$=-\frac{c}{48\pi}\int d^{2}\xi\sqrt{g}\left(\frac{1}{2}g^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi+R\phi\right),$ (3.10)
which is proportional to the Liouville action.
The same result can be obtained as follows. The renormalized partition function is written by the bare partition function with counterterm
$Z[g]=\lim_{\epsilon\to 0}Z[g_{0}]e^{\frac{c}{48\pi\epsilon_{2}}\int d^{2}\xi\sqrt{g}\,R},$ (3.11)
where $\epsilon_{2}=(2-D)/2$ and $c$ is a constant. The Liouville action arises from the counterterm as the finite part of the counterterm. Consider the Weyl transformation
$g_{\mu\nu}=e^{\phi}{\hat{g}}_{\mu\nu}.$ (3.12)
From the formula (A.2) given in Appendix A, the counterterm in (3.11) transforms into
$-\frac{c}{48\pi\epsilon_{2}}\int d^{D}x\sqrt{{\hat{g}}}\,e^{(D-2)\phi/2}\left(\hat{R}-(D-1)\hat{\square}\phi-\frac{(D-1)(D-2)}{4}(\partial_{\mu}\phi)^{2}\right),$ (3.13)
where the D’Lambertian and covariant derivatives on the rhs are made of the hatted metric ${\hat{g}}_{\mu\nu}$. Similar notation should be understood in what follows. The simple pole in the limit of $D\to 2\ (\epsilon_{2}\to 0)$ is the genuine counterterm. There are terms remaining finite in this limit. We can find these terms by expanding the integrand in (3.13) in (D2)(D-2) and keep the linear term. This gives
$\frac{c}{48\pi}\int d^{2}\xi\sqrt{{\hat{g}}}\left[\phi\big(\hat{R}-\hat{\square}\phi\big)-2\hat{\square}\phi-\frac{1}{2}(\partial_{\mu}\phi)^{2}\right]$
$=\frac{c}{48\pi}\int d^{2}\xi\sqrt{{\hat{g}}}\left[\frac{1}{2}(\partial_{\mu}\phi)^{2}+\hat{R}\phi\right],$ (3.14)
after partial integration, in agreement with Eq. (3.10).
The contribution of this Liouville action to the anomaly is given by
$c_{L}=1-c.$ (3.15)
Here 1 is the quantum contribution whereas c-c is the classical contribution coming from the shift of $\phi$ under the conformal transformation. The absence of the total conformal anomaly determines $c$ [14, 13]. To show it more explicitly, let us write the contribution to the central charge from the matter as $c^{\rm m}$. The contribution from the FP determinant or ghosts is 26-26, and that from the above Liouville action is 1c1-c
Therefore, to have a consistent theory, these must add up to zero:
$1-c-26+c^{\rm m}=0.$ (3.16)
This assures the consistency of the theory. That is, the theory remains unchanged even if $\hat{g}$ is replaced with $\hat{g}e^{\sigma}$.
Let us confirm the difference between critical and noncritical strings here. In critical strings, the total central charge of the matter field and gravity is zero, and the conformal modes do not appear in either the classical action or the quantum effective action. Therefore, we can simply forget about the conformal modes by regarding them as gauge degrees of freedom. On the other hand, in noncritical strings, the conformal modes do not appear in the classical action but do appear in the quantum effective action. The condition for consistently quantizing this effective action is (3.16).
We will apply this method to the four-dimensional conformal gravity to identify the four-dimensional Liouville theory.
4 Four-dimensional conformal anomaly and noncritical conformal gravity
4.1 Integrating trace anomaly
Using the method described in Sect. 3, here we derive the four-dimensional Liouvill action. Let us write
$\langle T^{\mu}{}_{\mu}\rangle=\frac{1}{(4\pi)^{2}}\left[c\,C_{\mu\nu\lambda\phi}^{2}-aR_{\rm GB}^{2}+b\square R\right].$ (4.1)
We consider the dimensional regularization and try to calculate the anomaly terms under the conformal transformation
$g_{\mu\nu}=e^{\phi}{\hat{g}}_{\mu\nu}.$ (4.2)
The Weyl tensor squared in $D$ dimensions is
$C_{\mu\nu\lambda\sigma}^{2}=R_{\mu\nu\lambda\sigma}^{2}-\frac{4}{D-2}R_{\mu\nu}^{2}+\frac{2}{(D-1)(D-2)}R^{2},$ (4.3)
and, according to the formulae in Appendix A, it transforms as
$\sqrt{g}\,C_{\mu\nu\lambda\sigma}^{2}\to\sqrt{{\hat{g}}}\,e^{((D-4)\phi/2}\hat{C}_{\mu\nu\lambda\rho}^{2}.$ (4.4)
Hence the counterterm
$\frac{c}{(4\pi)^{2}\epsilon}\int d^{D}x\sqrt{g}\,C_{\mu\nu\lambda\sigma}^{2},$ (4.5)
where $\epsilon=(4-D)/2$, produces finite terms, in the limit of $D\to 4$, as
$-\frac{c}{(4\pi)^{2}}\int d^{D}x\sqrt{{\hat{g}}}\,\phi\,\hat{C}_{\mu\nu\lambda\sigma}^{2}.$ (4.6)
Similarly the GB term transforms as
$\sqrt{g}R_{\rm GB}^{2}$ $=$ $\sqrt{{\hat{g}}}\,e^{(D-4)\phi/2}\Big[\hat{R}_{\rm GB}^{2}+2(D-3)\hat{R}^{\mu\nu}(2\hat{\nabla}_{\mu}\hat{\nabla}_{\nu}\phi-\partial_{\mu}\phi\partial_{\nu}\phi)-2(D-3)\hat{R}\,\hat{\square}\phi$
$-\frac{(D-3)(D-4)}{2}\hat{R}\,(\partial_{\mu}\phi)^{2}-(D-2)(D-3)(\hat{\nabla}_{\mu}\hat{\nabla}_{\nu}\phi)^{2}$
$+(D-2)(D-3)(\hat{\square}\phi)^{2}+(D-2)(D-3)(\hat{\nabla}_{\mu}\hat{\nabla}_{\nu}\sigma)\,\partial^{\mu}\phi\,\partial^{\nu}\phi$
$+\frac{(D-2)(D-3)^{2}}{2}\hat{\square}\phi(\partial_{\mu}\phi)^{2}+\frac{(D-1)(D-2)(D-3)(D-4)}{16}(\partial_{\mu}\phi)^{2}(\partial_{\nu}\phi)^{2}\Big].$
It is easy to make partial integration to get
$\int d^{D}x\sqrt{g}R_{\rm GB}^{2}\hskip-5.69054pt$ $=$ $\hskip-5.69054pt\int d^{D}x\sqrt{{\hat{g}}}\,e^{(D-4)\phi/2}\Big[\hat{R}_{\rm GB}^{2}-(D-3)(D-4)\hat{G}^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi$ (4.8)
$+\frac{(D-2)(D-3)(D-4)}{2}(\hat{\nabla}^{\mu}\hat{\nabla}^{\nu}\phi)\partial_{\mu}\phi\partial_{\nu}\phi$
$+\frac{(D-2)(D-3)(D-4)(D-5)}{16}\{(\partial_{\mu}\phi)^{2}\}^{2}\Big],$
where
$\hat{G}^{\mu\nu}=\hat{R}^{\mu\nu}-\frac{1}{2}\hat{R}{\hat{g}}^{\mu\nu},$ (4.9)
is the Einstein tensor. The terms depending on $\phi$ in Eq. (4.8) vanish for $D=4$, and this is consistent with the fact that GB terms are total derivatives in $D=4$. Hence the counterterm
$-\frac{a}{(4\pi)^{2}\epsilon}\int d^{D}x\sqrt{g}\,R_{GB}^{2},$ (4.10)
gives
$\frac{a}{(4\pi)^{2}}\int d^{4}x\Big[\phi\,\hat{R}_{\rm GB}^{2}-2\hat{G}^{\mu\nu}\hat{\nabla}_{\mu}\phi\hat{\nabla}_{\nu}\phi+2(\hat{\nabla}^{\mu}\hat{\nabla}^{\nu}\phi)\partial_{\mu}\phi\partial_{\nu}\phi-\frac{1}{4}\{(\partial_{\mu}\phi)^{2}\}^{2}\Big]$
$=\frac{a}{(4\pi)^{2}}\int d^{4}x\sqrt{g}\Big[\phi\,\hat{R}_{\rm GB}^{2}-2\hat{G}^{\mu\nu}\hat{\nabla}_{\mu}\phi\hat{\nabla}_{\nu}\phi-(\partial_{\mu}\phi)^{2}\hat{\square}\phi-\frac{1}{4}\{(\partial_{\mu}\phi)^{2}\}^{2}\Big].$ (4.11)
Collecting, our four-dimensional Liouville action would be
$S_{\rm eff}\hskip-5.69054pt$ $=$ $\hskip-5.69054pt\int d^{4}x\sqrt{{\hat{g}}}\,\Big[\frac{-c}{(4\pi)^{2}}\phi\,\hat{C}_{\mu\nu\lambda\sigma}^{2}+\frac{a}{(4\pi)^{2}}\Big(\phi\,\hat{R}_{\rm GB}^{2}-2\hat{G}^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi-(\hat{\square}\phi)(\partial_{\mu}\phi)^{2}-\frac{1}{4}\{(\partial_{\mu}\phi)^{2}\}^{2}\Big)\Big].$
Unfortunately this action contains nonlinear terms in the Liouville modes $\phi$ and is quite different from the two-dimensional case. Actually we should more carefully consider the quantization of the conformal gravity. We will see that this leads to better action as discussed in the following.
4.2 Four-dimensional noncritical conformal gravity
We consider again the conformal gravity (1.3) and write the partition function
$Z=\int[{\cal D}g]Z_{M}[g]e^{-\int d^{4}x\sqrt{g}(\frac{1}{2\lambda}C_{\mu\nu\lambda\rho}^{2}-\frac{1}{\rho}R_{\rm GB}^{2})}.$ (4.13)
To quantize this theory, we should first gauge fix the conformal invariance. For this purpose, we choose the conformal gauge fixing function (2.7). As discussed in Sect. 2, this introduces the partial gauge fixing and ghost terms (2.9). The partition function takes the form
$Z=\int[{\cal D}\!B{\cal D}{\hat{g}}{\cal D}\varphi{\cal D}b{\cal D}c]Z_{M}[{\hat{g}}e^{\varphi}]e^{-\int d^{4}x\sqrt{g}(\frac{1}{2\lambda}C_{\mu\nu\lambda\rho}^{2}-\frac{1}{\rho}R_{\rm GB}^{2}+{\cal L}_{\rm GF+FP})},$ (4.14)
where
${\cal L}_{\rm GF+FP}=\frac{1}{2\alpha}R^{2}+b\square c,$ (4.15)
from Eq. (2.9), where we have integrated out the auxiliary field $B$. We then gauge fix the diffeomorphism and perform the path integral by using the background formalism:
$Z=\int[{\cal D}\varphi]Z_{M}[{\hat{g}}]Z_{\rm G}[{\hat{g}}]e^{-S_{\rm eff}[{\hat{g}},\varphi]},$ (4.16)
where $Z_{\rm G}[{\hat{g}}$] is the gravity part for the background metric $^{\hat{g}}$. In this formulation, any choice of $2\alpha=\xi$ is allowed since it is a gauge parameter. We could also consider the contribution of the finite $R^{2}$ term to the Liouville action. We will see that the consistency of the anomaly uniquely singles out its coefficient. With this value, we find that the Liouville mode becomes free, while other choice gives nontrivial interacting theory.
Let us consider a combination of the finite term (4.10) from the counterterm $R_{GB}^{2}$ and the finite term $R^{2}$:
$\left.-\frac{a}{(4\pi)^{2}\epsilon}\int d^{D}x\sqrt{g}R_{GB}^{2}\right|_{\rm finite}+\frac{e}{(4\pi)^{2}}\int d^{D}x\sqrt{g}R^{2},$ (4.17)
where we have chosen the coefficient of the first term such that it has the correct normalization to give the corresponding term in (LABEL:4dl), and have written the coefficient of the second term as $\frac{e}{(4\pi)^{2}}$ for convenience. We use the method of Sect. 3 to extract the finite term as a Liouville action. We would like to determine the coefficient $e$ by imposing the consistency discussed in Appendix C. Under the conformal transformation $g_{\mu\nu}\to e^{\phi}{\hat{g}}_{\mu\nu}$, we find from Eqs. (4.17) and (A.5):
$\int d^{4}x\sqrt{{\hat{g}}}\,\frac{1}{(4\pi)^{2}}\Big[$ $\hskip-17.07164pta\Big\{\phi\,\hat{R}_{\rm GB}^{2}-2\hat{G}^{\mu\nu}\hat{\nabla}_{\mu}\phi\hat{\nabla}_{\nu}\phi-(\partial_{\mu}\phi)^{2}\hat{\square}\phi-\frac{1}{4}\{(\partial_{\mu}\phi)^{2}\}^{2}\Big\}$
$\hskip-17.07164pt+e\Big\{\hat{R}^{2}-6\hat{R}\hat{\square}\phi-3\hat{R}(\partial_{\mu}\phi)^{2}+9(\hat{\square}\phi)^{2}+9\hat{\square}\phi(\partial_{\mu}\phi)^{2}+\frac{9}{4}\{(\partial_{\mu}\phi)^{2}\}^{2}\Big\}\Big]$
$=\int d^{4}e\sqrt{{\hat{g}}}\,\frac{1}{(4\pi)^{2}}\Big[$ $\hskip-17.07164pt\Big\{a\,\phi\,\hat{R}_{\rm GB}^{2}+e\hat{R}^{2}-6e\phi\hat{\square}\hat{R}-2a\hat{G}^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi-3e\hat{R}(\partial_{\mu}\phi)^{2}+9e(\hat{\square}\phi)^{2}$ (4.18)
$-(a-9e)(\partial_{\mu}\phi)^{2}\hat{\square}\phi-\frac{(a-9e)}{4}\{(\partial_{\mu}\phi)^{2}\}^{2}\Big\}.$
The conformal anomaly at the one-loop level gets contribution only from the quadratic terms in $\phi$. From Eq. (4.18), they are given as
$\int d^{4}x\sqrt{{\hat{g}}}\,\frac{9e}{(4\pi)^{2}}\Big[(\hat{\square}\phi)^{2}-\frac{1}{3}\hat{R}(\partial_{\mu}\phi)^{2}-\frac{2a}{9e}\hat{G}^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi\Big]$
$=\int d^{4}x\sqrt{{\hat{g}}}\,\frac{9e}{(4\pi)^{2}}\phi\Big[\hat{\square}^{2}+\Big\{x\hat{R}^{\mu\nu}+y\hat{R}{\hat{g}}^{\mu\nu}\Big\}\hat{\nabla}_{\mu}\hat{\nabla}_{\nu}+\frac{1}{3}(\partial_{\mu}\hat{R})\partial^{\mu}\Big]\phi,$ (4.19)
Here in going from the first line to the next, we have made the partial integration, used the Bianchi identity and set
$x=\frac{2a}{9e},\qquad y=\frac{1}{3}-\frac{x}{2}.$ (4.20)
Using the formula from [15], we can calculate their contribution to the anomaly:
$\hskip-22.76219pt\frac{1}{(4\pi)^{2}}\frac{\Gamma(2)}{2\Gamma(3)}\left[2\Big(\frac{1}{90}\hat{R}_{\mu\nu\rho\phi}^{2}-\frac{1}{90}\hat{R}_{\mu\nu}^{2}+\frac{1}{36}\hat{R}^{2}+\frac{1}{15}\hat{\square}\hat{R}\Big)+\frac{2}{9}(x+4y)\hat{\square}\hat{R}-\frac{5}{18}(x+2y)\hat{\square}\hat{R}\right$.
$\hskip-14.22636pt\left.+\frac{1}{24}(x+4y)^{2}\hat{R}^{2}+\frac{1}{12}(x^{2}\hat{R}_{\mu\nu}^{2}+(2xy+4y^{2})\hat{R}^{2})+\frac{1}{6}(x+4y)\hat{R}^{2}-\frac{1}{3}(x\hat{R}_{\mu\nu}^{2}+y\hat{R}^{2})+\frac{1}{3}\hat{\square}\hat{R}\right$],
which may be rewritten as
$\frac{1}{1440(4\pi)^{2}}$ $\hskip-14.22636pt\left[3(5x^{2}-20x+4)\hat{C}_{\mu\nu\rho\phi}^{2}-(15x^{2}-60x+4)\hat{R}_{GB}^{2}+25(x-2)^{2}\hat{R}^{2}\right$. (4.22)
$\left.+16(13-5x)\hat{\square}\hat{R}\right$].
Requiring that there should be no $R^{2}$ term, as follows from the consistency condition discussed in Appendix C, we find uniquely
$x=2,\qquad y=-\frac{2}{3},$ (4.23)
which in turn gives the condition
$e=\frac{a}{9}.$ (4.24)
The action (4.19) reduces to
$\int d^{4}x\sqrt{\hat{g}}\,\frac{a}{(4\pi)^{2}}\phi\Big[\hat{\square}^{2}+\Big\{2\hat{R}^{\mu\nu}-\frac{2}{3}\hat{R}\hat{g}^{\mu\nu}\Big\}\hat{\nabla}_{\mu}\hat{\nabla}_{\nu}+\frac{1}{3}(\partial_{\mu}\hat{R})\partial^{\mu}\Big]\phi$
$=\int d^{4}x\sqrt{\hat{g}}\,\frac{a}{(4\pi)^{2}}\Big[(\hat{\square}\phi)^{2}-2\hat{R}^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi+\frac{2}{3}\hat{R}(\partial_{\mu}\phi)^{2}\Big],$ (4.25)
and it is intriguing that for this choice the interaction terms in $\phi$ in Eq. (4.18) drop out and the theory is free (up to gravity interaction). This action is precisely the conformally invariant one for a scalar field, as we see in eq. (B.7) in Appendix B. So our Liouville action is expressed as
$S_{4D,L}[g,\phi]=\lim_{D\to 4}\left(S_{c}[ge^{\phi}]-S_{c}[g]\right),$ (4.26)
where
$S_{c}[g]=\int d^{D}x\sqrt{g}\ \frac{1}{(4\pi)^{2}}\Big(\frac{c}{\epsilon}\,C_{\mu\nu\rho\sigma}^{2}-\frac{a}{\epsilon}R_{GB}^{2}+\frac{a}{9}R^{2}\Big).$ (4.27)
Collecting all terms, our anomaly or Liouville action is
$S_{4D,L}[{\hat{g}},\phi]\hskip-5.69054pt$ $=$ $\hskip-5.69054pt\int d^{4}x\sqrt{{\hat{g}}}\ \frac{1}{(4\pi)^{2}}\Big[-c\phi C_{\mu\nu\rho\sigma}^{2}+a\Big(\phi R_{GB}^{2}-2\hat{G}^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi+(\hat{\square}\phi)^{2}$ (4.28)
$\hskip 85.35826pt-\frac{2}{3}\phi\hat{\square}\hat{R}-\frac{1}{3}\hat{R}(\partial_{\mu}\phi)^{2}\Big)\Big].$
It is quite remarkable that the requirement that we should have consistent conformal anomaly leads to the result that the Liouville mode becomes free field (up to gravity interactions), as in two-dimensional Liouville theory. [^3] We can also confirm that this can be achieved if we use $R_{\mu\nu}^{2}$ for the local term but factor 3 difference. From the formula (A.4), we have $\hskip-28.45274pt\int d^{4}x\sqrt{g}\,R_{\mu\nu}{}^{2}=\int d^{4}x\sqrt{{\hat{g}}}\Big[\hat{R}_{\mu\nu}{}^{2}-2\hat{R}_{\mu\nu}\hat{\nabla}^{\mu}\hat{\nabla}^{\nu}\phi+\hat{R}_{\mu\nu}\partial^{\mu}\phi\partial^{\nu}\phi-\hat{R}\hat{\square}\phi-\hat{R}(\partial_{\mu}\phi)^{2}$ $\hskip 28.45274pt+(\hat{\nabla}_{\mu}\hat{\nabla}_{\nu}\phi)^{2}+2(\hat{\square}\phi)^{2}-(\hat{\nabla}_{\mu}\hat{\nabla}_{\nu}\phi)\partial^{\mu}\phi\partial^{\nu}\phi+\frac{5}{2}\hat{\square}\phi(\partial_{\mu}\phi)^{2}+\frac{3}{4}(\partial_{\mu}\phi)^{2}(\partial_{\nu}\phi)^{2}\Big]$ $\hskip-14.22636pt=\int d^{4}x\sqrt{{\hat{g}}}\Big[\hat{R}_{\mu\nu}{}^{2}-2\phi\hat{\square}\hat{R}+3(\hat{\square}\phi)^{2}-\hat{R}(\partial_{\mu}\phi)^{2}+3\hat{\square}\phi(\partial_{\mu}\phi)^{2}+\frac{3}{4}(\partial_{\mu}\phi)^{2}(\partial_{\nu}\phi)^{2}\Big].$ (4.29) It is then easy to show that the Liouville mode becomes free and we can obtain consistent anomaly if we add this term with the factor $e/3$. One can also check using (A.3) that $R_{\mu\nu\rho\sigma}^{2}$ term also gives the same terms as $R_{\mu\nu}^{2}$ and so we could use this. However $R^{2}$ is singled out since it can be introduced as a gauge fixing term. Other choice of $e$ is not allowed since it cannot give consistent trace anomaly, and furthermore it would give interacting Liouville mode.
Before beginning the quantization of Liouville theory, let us first confirm how the Liouville action (4.28) transforms under conformal transformations at the classical level. As is evident from its construction (4.26), we easily obtain the addition theorem
$S_{4D,L}[g,\phi_{1}+\phi_{2}]=S_{4D,L}[ge^{\phi}_{1},\phi_{2}]+S_{4D,L}[g,\phi_{1}].$ (4.30)
By substituting $\phi_{1}=\sigma,\phi_{2}=\phi-\sigma$ to this equation and moving terms properly, we find how the Liouville action transforms under the conformal transformation at the classical level:
$S_{4D,L}[ge^{\sigma},\phi-\sigma]-S_{4D,L}[g,\sigma]=-S_{4D,L}[g,\sigma].$ (4.31)
Here the conformal transformation is defined by
$g\mapsto ge^{\sigma},\phi\mapsto\phi-\sigma.$ (4.32)
As in the two-dimensional case (3.15), the anomaly coefficients of the Liouville field (4.28) are expressed as the sum of classical and quantum contributions.
$c_{4D,L}=-\frac{1}{30}-c^{\prime},$
$a_{4D,L}=-\frac{7}{180}-a^{\prime}.$ (4.33)
Here, $c^{\prime}$ and $a^{\prime}$ are classical contributions coming from (4.31), and the quantum one-loop anomaly from (4.25) is given in (4.22) with $x$ and $y$ in Eq. (4.23):
$\frac{1}{(4\pi)^{2}}\left[-\frac{1}{30}(\hat{C}_{\mu\nu\rho\sigma}^{2}-\hat{\square}\hat{R})+\frac{7}{180}\hat{R}_{GB}^{2}\right].$ (4.34)
With these preparations in place, we can discuss the quantization of noncritical conformal gravity. Our starting point is the following classically conformal theory:
$S[g,\varphi^{\text{m}}]=S^{\text{m}}[g,\varphi^{\text{m}}]+\int d^{4}x\sqrt{g}\frac{1}{\lambda}\,C_{\mu\nu\rho\sigma}^{2},$ (4.35)
where $S^{\text{m}}[g,\varphi^{\text{m}}]$ is the action for conformal matter fields $\varphi^{\text{m}}$ with anomaly coefficients $c^{\text{m}}$ and $a^{\text{m}}$. We consider the noncritical case, that is, where (2.17) is not satisfied. We then separate the conformal modes by decomposing $g$ as
$g_{\mu\nu}={\hat{g}}_{\mu\nu}e^{\phi}.$ (4.36)
Here the degrees of freedom of the conformal mode of $^{\hat{g}}$ are suppressed by the gauge fixing of its conformal symmetry. Thus (4.35) is equivalent to
$S[{\hat{g}},\phi,\varphi^{\text{m}}]=\Big(S^{\text{m}}[g,\varphi^{\text{m}}]+\int d^{4}x\sqrt{g}\frac{1}{\lambda}\,C_{\mu\nu\rho\sigma}^{2}\Big)_{g={\hat{g}}e^{\phi}}+\int d^{4}x\sqrt{{\hat{g}}}\Big(\frac{1}{2\alpha}\hat{R}^{2}+b\square_{{\hat{g}}}c\Big),$ (4.37)
If we ignore the conformal anomalies, the contributions to the partition function of the first term in (4.37) does not depend on $\phi$ because the action is conformally invariant. Taking the anomalies into account, or considering the counter terms of the form (4.27), we have
$S[{\hat{g}},\phi,\varphi^{\text{m}}]=S^{\text{m}}[{\hat{g}},\varphi^{\text{m}}]+\int d^{4}x\sqrt{{\hat{g}}}\frac{1}{\lambda}\,\hat{C}_{\mu\nu\rho\sigma}^{2}+S_{4D,L}[{\hat{g}},\phi]+\int d^{4}x\sqrt{{\hat{g}}}\Big(\frac{1}{2\alpha}\hat{R}^{2}+b\square_{{\hat{g}}}c\Big).$ (4.38)
As in the two-dimensional case, the Liouville action can be regarded as the Jacobian for the change of measure in the path integral:
$\mathcal{D}_{g}g\,\mathcal{D}_{g}\phi\,\mathcal{D}_{g}\varphi^{\text{m}}=\mathcal{D}_{\hat{g}}{\hat{g}}\,\mathcal{D}_{\hat{g}}\phi\,\mathcal{D}_{\hat{g}}\varphi^{\text{m}}\,e^{-S_{4D,L}[{\hat{g}},\phi]},$ (4.39)
where $\mathcal{D}_{g}\varphi$ stands for the path measure for the field $\varphi$ defined on the back ground metric $g_{\mu\nu}$. As with the two-dimensional case, the parameters appearing in the Liouville action are difficult to determine a priori because they undergo renormalization. However, they can be determined from the overall consistency as follows.
After rewriting $^{\hat{g}}$ to $g$ in (4.38), we have the total action:
$S_{\text{total}}[g,\phi,\varphi^{\text{m}}]=S^{\text{m}}[g,\varphi^{\text{m}}]+\int d^{4}x\sqrt{g}\Big(\frac{1}{\lambda}\,C_{\mu\nu\rho\sigma}^{2}+\frac{1}{2\alpha}\,R^{2}+b\square c\Big)+S_{4D,L}[g,\phi],$ (4.40)
where the last term $S_{4D,L}[g,\phi]$ contains two undetermined coefficients $a$ and $c$ (see (4.28)). In fact we can determine them by the requirement that $S_{\text{total}}$ have no anomaly for the BRS transformation of the conformal transformation, (2.2), (2.3), (2.5) and
$\delta_{B}\phi=-\delta\lambda\,c.$ (4.41)
This is nothing other than the entire system, including the Liouville field and the ghost fields for fixing the gauge of the conformal transformation, being invariant under the conformal transformation:
$c_{\rm total}=c^{\text{conf}}+c^{\text{m}}+c_{4D,L}=0,$
$a_{\rm total}=a^{\text{conf}}+a^{\text{m}}+a_{4D,L}=0.$ (4.42)
Substituting (2.14), (2.16), and (4.33), we have
$\frac{199}{30}+\frac{1}{120}(N_{S}+6N_{F}+12N_{V})-\frac{1}{30}-c^{\prime}=0,$
$\frac{87}{20}+\frac{1}{360}(N_{S}+11N_{F}+62N_{V})-\frac{7}{180}-a^{\prime}=0.$ (4.43)
If this condition is satisfied, the BRS transformation is anomaly free at least at the 1-loop level and possesses nilpotency. [^4] Note that $\alpha$ in (4.40) is arbitrary as long as BRS symmetry is anomaly free. On the other hand, when considering higher-order corrections, the beta functions become more complex, and it is not self-evident whether $c_{\rm total}$ and $a_{\rm total}$ can remain zero without receiving renormalization.
4.3 Comments on Riegert proposal
On the other hand, Riegert proposed the four-dimensional Liouville action as follows [10].
Recall that the scalar curvature transforms as (A.2) in Appendix A. We can calculate
$\sqrt{g}\,\square R\hskip-5.69054pt$ $=$ $\hskip-5.69054pt\partial_{\mu}(\sqrt{g}\,g^{\mu\nu}\partial_{\nu}R)$ (4.44)
$=$ $\hskip-5.69054pt\sqrt{{\hat{g}}}\,e^{(D-4)\phi/2}\Big[\hat{\square}\hat{R}-\frac{D-4}{2}\hat{R}(\partial_{\mu}\phi)^{2}-\hat{R}\,\hat{\square}\phi+\frac{D-6}{2}\partial^{\mu}\hat{R}\,\partial_{\mu}\phi$
$-\frac{(D-1)(D-2)}{2}\hat{R}^{\mu\nu}\partial_{\mu}\phi\,\partial_{\nu}\phi+(D-1)(\hat{\square}\phi)^{2}$
$-\frac{(D-1)(D-2)(D-6)}{4}(\hat{\nabla}^{\mu}\partial^{\nu}\phi)\partial_{\mu}\phi\partial_{\nu}\phi+\frac{(D-1)(3D-10)}{4}\hat{\square}\phi(\partial_{\mu}\phi)^{2}$
$+\frac{(D-1)(D-2)(D-4)}{8}(\partial_{\mu}\phi)^{2}(\partial_{\nu}\phi)^{2}-(D-1)(D-4)\partial_{\mu}\hat{\square}\phi\partial^{\mu}\phi$
$-(D-1)\hat{\square}^{2}\phi-\frac{(D-1)(D-2)}{2}(\hat{\nabla}_{\mu}\hat{\nabla}_{\nu}\phi)^{2}\Big].$
Together with the results in (A.3) –(A.5) or (LABEL:gbconforig), we get
$\sqrt{g}\Big[R_{\rm GB}^{2}-\frac{2}{3}\square R\Big]\hskip-5.69054pt$ $=$ $\hskip-5.69054pt\sqrt{{\hat{g}}}\,e^{(D-4)\phi/2}\Big[\hat{R}_{\rm GB}^{2}-\frac{2}{3}\hat{\square}\hat{R}+\frac{2(D-1)}{3}\hat{\square}^{2}\phi-\frac{1}{3}(D-6)\partial^{\mu}\hat{R}\partial_{\mu}\phi$ (4.45)
$\hskip-14.22636pt-\frac{2}{3}(3D-10)\hat{R}\hat{\square}\phi+4(D-3)\hat{R}^{\mu\nu}\hat{\nabla}_{\mu}\hat{\nabla}_{\nu}\phi-\frac{(D-4)(3D-11)}{6}\hat{R}(\partial_{\mu}\phi)^{2}$
$\hskip-14.22636pt+\frac{(D-4)(D-5)}{3}\hat{R}^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi-\frac{2}{3}(D-2)(D-4)(\hat{\nabla}_{\mu}\hat{\nabla}_{\nu}\phi)^{2}$
$\hskip-14.22636pt+\frac{(D-4)(3D-5)}{3}(\hat{\square}\phi)^{2}+\frac{(D-2)(D+3)(D-4)}{6}(\hat{\nabla}^{\mu}\hat{\nabla}^{\nu}\phi)\partial_{\mu}\phi\,\partial_{\nu}\phi$
$\hskip-14.22636pt+\frac{(D-1)(D-2)(D-4)(3D-13)}{48}\{(\partial_{\mu}\phi)^{2}\}^{2}+\frac{2}{3}(D-1)(D-4)\partial^{\mu}\hat{\square}\phi\partial_{\mu}\phi$
$\hskip-14.22636pt-\frac{1}{6}(D-4)(3D^{2}-15D+16)\hat{\square}\phi(\partial_{\mu}\phi)^{2}\Big].$
If we set $D=4$, there remains only linear terms in $\phi$.
$\sqrt{g}\Big[R_{\rm GB}^{2}-\frac{2}{3}\square R\Big]\hskip-5.69054pt$ $=$ $\hskip-5.69054pt\sqrt{{\hat{g}}}\Big[\hat{R}_{\rm GB}^{2}-\frac{2}{3}\hat{\square}\hat{R}+2\hat{\square}^{2}\phi+\frac{2}{3}\partial^{\mu}\hat{R}\partial_{\mu}\phi-\frac{4}{3}\hat{R}\hat{\square}\phi+4\hat{R}^{\mu\nu}\hat{\nabla}_{\mu}\hat{\nabla}_{\nu}\phi\Big].\qquad\quad$ (4.46)
It is an interesting observation that there remain only linear terms in $\phi$ in this combination [10]. It was then proposed that these terms multiplied by $\phi/2$ give the anomaly action and is widely used since then [16]. However this does not follow the procedure to integrate the trace anomaly and it is not clear why this has anything to do with the four-dimensional Liouville theory. Moreover these terms are total derivatives, and there is no rationale why we can get the anomaly action just by multiplying $\phi$ to these terms. On the other hand, it is clear that (4.28) gives the anomaly under the shift of the conformal mode $\phi$ in the first order in $\phi$.
According to the general wisdom from two-dimensional quantum gravity, we should rather consider the counterterm
$-\frac{1}{\epsilon}\int d^{D}x\sqrt{g}\big[R_{\rm GB}^{2}-\frac{2}{3}\square R\big],$ (4.47)
and make the Weyl transformation $g_{\mu\nu}\to e^{\phi}{\hat{g}}_{\mu\nu}$ to extract the finite term in the limit of $D\to 4$. This produces
$\int d^{4}x$ $\hskip-17.07164pt\sqrt{{\hat{g}}}\Big[\phi\Big\{\hat{R}_{\rm GB}^{2}-\frac{2}{3}\hat{\square}\hat{R}+2\hat{\square}^{2}\phi+\frac{2}{3}\partial^{\mu}\hat{R}\,\partial_{\mu}\phi-\frac{4}{3}\hat{R}\hat{\square}\phi+4\hat{R}^{\mu\nu}\hat{\nabla}_{\mu}\hat{\nabla}_{\nu}\phi\Big\}+\frac{4}{3}\hat{\square}^{2}\phi-\frac{2}{3}\partial^{\mu}\hat{R}\partial_{\mu}\phi$ (4.48)
$-4\hat{R}\hat{\square}\phi+8\hat{R}^{\mu\nu}\hat{\nabla}_{\mu}\hat{\nabla}_{\nu}\phi-\frac{1}{3}\hat{R}(\partial_{\mu}\phi)^{2}-\frac{2}{3}\hat{R}^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi-\frac{8}{3}(\hat{\nabla}_{\mu}\hat{\nabla}_{\nu}\phi)^{2}+\frac{14}{3}(\hat{\square}\phi)^{2}$
$+\frac{14}{3}(\hat{\nabla}^{\mu}\partial^{\nu}\phi)\partial_{\mu}\phi\,\partial_{\nu}\phi-\frac{1}{4}\{(\partial_{\mu}\phi)^{2}\}^{2}-4(\hat{\square}\phi)^{2}-\frac{4}{3}\hat{\square}\phi(\partial_{\mu}\phi)^{2}\Big].\qquad$
The above proposal is just to keep the curly bracket terms, but this overlooks the remaining terms. It is not clear how this can be justified. Indeed it is easy to check that (4.48) reduces to (4.11) upon partial integration. This is to be expected because the additional term in (4.45) is a total derivative.
5 Summary
In this paper, we have first recalled the puzzle associated the beta functions for general quadratic curvature theory and the conformal gravity, and show that the puzzle is resolved by considering the partial gauge fixing of the conformal gauge symmetry with the BRST transformation. Extending the formulation with the BRST symmetry to the noncrtical conformal gravity, we have proposed that the four-dimensional Liouville theory should be given by Eq. (4.28), which is derived by deliberately adding the finite $R^{2}$ term (4.17) such that the resulting action satisfies the requirement of the consistency of the conformal anomaly and becomes a free quadratic action in the conformal mode, similar to two-dimensional Liuoville theory. We have also given the condition that the BRST symmetry is anomaly free, such that the quantum noncritical conformal gravity is consistent. Finally, we have contrasted this approach with the widely-used Riegert proposal [10] for the four-dimensional Liouville action, showing that our formulation is a more robust, consistent one derived through the quantization procedure.
Acknowledgments
H.K. is partially supported by the Ministry of Science and Technology, R.O.C. (MOST 111-2811-M-002-016), and by National Taiwan University. H.K. also thanks Prof. Shin-Nan Yang and his family for their kind support through the Chin-Yu chair professorship. N.O. would like to thank Ningbo University, where this work was completed, for their kind hospitality and support.
Appendix A Conformal transformation of curvature terms
If we make the Weyl transformation
$g_{\mu\nu}=e^{\sigma}\tilde{g}_{\mu\nu},$ (A.1)
the curvature tensors in $D$ dimensions transforms as
$R_{\mu\nu}=\tilde{R}_{\mu\nu}-\frac{D-2}{2}\tilde{\nabla}_{\mu}\tilde{\nabla}_{\nu}\sigma-\frac{1}{2}\tilde{g}_{\mu\nu}\tilde{\square}\sigma+\frac{D-2}{4}[\partial_{\mu}\sigma\partial_{\nu}\sigma-\tilde{g}_{\mu\nu}(\partial_{\alpha}\sigma)^{2}],$
$R=e^{-\sigma}\Big[\tilde{R}-(D-1)\tilde{\square}\sigma-\frac{(D-1)(D-2)}{4}(\partial_{\mu}\sigma)^{2}\Big].$ (A.2)
A tilde indicates that the quantity is evaluated on the tilded metric $\tilde{g}$ and the indices are raised, lowered and contracted by it.
The curvature square transforms as
$\sqrt{g}\,R^{\mu}{}_{\nu\alpha\beta}{}^{2}\!\!$ $=$ $\!\!\sqrt{\tilde{g}}\,e^{(D-4)\sigma/2}\Big[\tilde{R}^{\mu}{}_{\nu\alpha\beta}{}^{2}-4\tilde{R}_{\mu\nu}\tilde{\nabla}^{\mu}\tilde{\nabla}^{\nu}\sigma+2\tilde{R}_{\mu\nu}\partial^{\mu}\sigma\partial^{\nu}\sigma-\tilde{R}(\partial_{\mu}\sigma)^{2}$ (A.3)
$+(D-2)(\tilde{\nabla}_{\mu}\tilde{\nabla}_{\nu}\sigma)^{2}+(\tilde{\square}\sigma)^{2}-(D-2)(\tilde{\nabla}_{\mu}\tilde{\nabla}_{\nu}\sigma)\partial^{\mu}\sigma\partial^{\nu}\sigma+(D-2)\tilde{\square}\sigma(\partial_{\nu}\sigma)^{2}$
$+\frac{(D-1)(D-2)}{8}\{(\partial_{\mu}\sigma)^{2}\}^{2}\Big],$
$\sqrt{g}\,R_{\mu\nu}{}^{2}\!\!$ $=$ $\!\!\sqrt{\tilde{g}}\,e^{(D-4)\sigma/2}\Big[\tilde{R}_{\mu\nu}{}^{2}-(D-2)\tilde{R}_{\mu\nu}\tilde{\nabla}^{\mu}\tilde{\nabla}^{\nu}\sigma-\tilde{R}\tilde{\square}\sigma+\frac{D-2}{2}\tilde{R}_{\mu\nu}\partial^{\mu}\sigma\partial^{\nu}\sigma$ (A.4)
$-\frac{D-2}{2}\tilde{R}(\partial_{\mu}\sigma)^{2}+\frac{(D-2)^{2}}{4}(\tilde{\nabla}_{\mu}\tilde{\nabla}_{\nu}\sigma)^{2}+\frac{3D-4}{4}(\tilde{\square}\sigma)^{2}$
$-\frac{(D-2)^{2}}{4}(\tilde{\nabla}_{\mu}\tilde{\nabla}_{\nu}\sigma)\partial^{\mu}\sigma\partial^{\nu}\sigma+\frac{(D-2)(2D-3)}{4}\tilde{\square}\sigma(\partial_{\mu}\sigma)^{2}$
$+\frac{(D-1)(D-2)^{2}}{16}(\partial_{\mu}\sigma)^{2}(\partial_{\nu}\sigma)^{2}\Big],$
$\sqrt{g}\,R^{2}\!\!$ $=$ $\!\!\sqrt{\tilde{g}}\,e^{(D-4)\sigma/2}\Big[\tilde{R}^{2}-2(D-1)\tilde{R}\tilde{\square}\sigma-\frac{(D-1)(D-2)}{2}\tilde{R}(\partial_{\mu}\sigma)^{2}+(D-1)^{2}(\tilde{\square}\sigma)^{2}$ (A.5)
$+\frac{(D-1)^{2}(D-2)}{2}\tilde{\square}\sigma(\partial_{\nu}\sigma)^{2}+\frac{(D-1)^{2}(D-2)^{2}}{16}\{(\partial_{\mu}\sigma)^{2}\}^{2}\Big],$
The covariant derivatives on the rhs are all constructed by tilded metric $\tilde{g}_{\mu\nu}$.
Appendix B Conformally invariant scalar fields
Consider the usual kinetic term for a scalar field
$\int d^{4}x\sqrt{g}\ \frac{1}{2}g^{\mu\nu}\partial_{\mu}\varphi\partial_{\nu}\varphi.$ (B.1)
We consider the conformal transformation
$g_{\mu\nu}=e^{\sigma}\tilde{g}_{\mu\nu}.$ (B.2)
To make the action (B.1) invariant under this transformation, we should transform the scalar field as
$\varphi=e^{-\sigma/2}\tilde{\varphi}$. (B.3)
Using the transformation property (A.2), we then find that the combination
$\int d^{4}x\sqrt{g}\ \frac{1}{2}\left[g^{\mu\nu}\partial_{\mu}\varphi\partial_{\nu}\varphi+\frac{1}{6}R\varphi^{2}\right],$ (B.4)
is invariant under the conformal transformation.
Let us consider conformally invariant theory with higher derivative
$\int d^{4}x\sqrt{g}\ \frac{1}{2}(\square\varphi)^{2}.$ (B.5)
To make this invariant under the transformation, we must transform the scalar field as
$\varphi=\tilde{\varphi}$. (B.6)
Again using the transformation property (A.2), we find that the combination
$\int d^{4}x\sqrt{g}\ \frac{1}{2}\left[(\square\varphi)^{2}-2R^{\mu\nu}\partial_{\mu}\varphi\partial_{\nu}\varphi+\frac{2}{3}Rg^{\mu\nu}\partial_{\mu}\varphi\partial_{\nu}\varphi\right],$ (B.7)
is invariant under the conformal transformation.
On the other hand, if we just take the d’Alembertian operator square $\square^{2}$, according to the formula in Ref. [15], it gives the anomaly
$\frac{1}{(4\pi)^{2}}\left[\frac{1}{120}\tilde{C}_{\mu\nu\rho\sigma}^{2}+\frac{1}{30}\tilde{\square}\tilde{R}-\frac{1}{360}\tilde{R}_{GB}^{2}+\frac{1}{72}\tilde{R}^{2}\right].$ (B.8)
The coefficients of $\tilde{C}_{\mu\nu\rho\sigma}^{2}$ and $\tilde{R}_{GB}^{2}$ just correspond to those of two scalar matter fields. If we consider only normal kinetic term -$\square$, it gives
$\frac{1}{(4\pi)^{2}}\left[\frac{1}{120}\tilde{C}_{\mu\nu\rho\sigma}^{2}+\frac{1}{30}\tilde{\square}\tilde{R}-\frac{1}{360}\tilde{R}_{GB}^{2}+\frac{1}{72}\tilde{R}^{2}\right].$ (B.9)
Surprizing enough, this also gives the same contribution as $\square^{2}$. For the conformally invariant operator $\square+\frac{1}{6}R$, it gives
$\frac{1}{(4\pi)^{2}}\left[\frac{1}{120}\tilde{C}_{\mu\nu\rho\sigma}^{2}+\frac{1}{180}\tilde{\square}\tilde{R}-\frac{1}{360}\tilde{R}_{GB}^{2}\right].$ (B.10)
Appendix C Consistency of the trace anomaly
The general form of the quantum trace anomaly in four dimensions is [17]
$\langle T_{\mu}^{\mu}\rangle$ $=$ $-\frac{2}{\sqrt{g}}g_{\mu\nu}\frac{\delta S_{\rm eff}}{\delta g_{\mu\nu}}(g_{\mu\nu}=e^{\phi}{\hat{g}}_{\mu\nu})$ (C.1)
$=$ $-\frac{\delta S_{\rm eff}}{\delta{\phi}}$
$\sim$ $R^{2}+AR_{\mu\nu}R^{\mu\nu}+BR_{\mu\nu\lambda\rho}R^{\mu\nu\lambda\rho}.$
up to overall factor. The consistency of the trace anomaly requires that
$\delta^{2}S_{\rm eff}=0.$ (C.2)
From the transformation property given in Appendix A, we see under infinitesimal transformation $\delta\sigma$, in four dimensions,
$\delta^{2}S_{\rm eff}=\int d^{4}x\sqrt{g}\Big[-6R\,\square\delta\phi-A(2R^{\mu\nu}\,\nabla_{\mu}\nabla_{\nu}\delta\phi+R\,\square\delta\phi)-4BR^{\mu\nu}\nabla_{\mu}\nabla_{\nu}\delta\phi\Big]\wedge\delta\phi.$ (C.3)
Here and in what follows, a tilde indicates that the quantity is evaluated on the hatted metric $^{\hat{g}}$ and the indices are raised, lowered and contracted by it. Making the partial integration to those terms with Ricci tensor, and using the Bianchi identity
$\nabla_{\mu}R^{\mu\nu}=\frac{1}{2}\nabla^{\nu}R,$ (C.4)
and making back the partial integration, (C.3) gives
$\delta^{2}S_{\rm eff}=-\int d^{4}\sqrt{g}\,2(3+A+B)R\,\square\delta\phi\wedge\delta\phi\Big$. (C.5)
Namely the consistency gives
$A+B+3=0.$ (C.6)
Two independent solutions are
$A=-4,\qquad B=1,$ (C.7)
which is precisely the GB term (1.5), and
$A=-6,\qquad B=3$ (C.8)
which corresponds to the Weyl tensor square (1.4). To summarize, the trace anomaly is restricted to GB term and Weyl tensor squared by the consistency. The important point is that there is no $R^{2}$ term.
The partition function is given by
$Z[ge^{\phi}]=Z[g]e^{S_{\rm eff}[\phi,g]}.$ (C.9)
When the consistency condition is satisfied, we have
$Z[ge^{\phi}e^{\phi^{\prime}}]$ $=$ $Z[ge^{\phi}]e^{-S_{\rm eff}[\phi^{\prime},ge^{\phi}]}$ (C.10)
$=$ $Z[g]e^{-S_{\rm eff}[\phi,g]-S_{\rm eff}[\phi^{\prime},ge^{\phi}]}$
$=$ $Z[g]e^{-S_{\rm eff}[\phi+\phi^{\prime},g]}.$
This means that
$S_{\rm eff}[\phi+\phi^{\prime},g]=S_{\rm eff}[\phi,g]+S_{\rm eff}[\phi^{\prime},ge^{\phi}].$ (C.11)